how can we find the l graph if certain result is given without actual plotting say s=t^2 or s=T^3

Grade:11

2 Answers

vikas askiitian expert
13 years ago

if  s=t3       .........1

  v=ds/dt=3t2..................2

 a=dv/dt=6t....................3

solving 2 and 3 

 a2 = 12v

it means graph bw accleration and velocity is parabola...

in same way graph of v  vs  s  or  a  vs  s can be made....

Aman Bansal
13 years ago

dear arpit,
if you want to plot a graph then you can use calculus without using the values

for example if you want to plot y=x^2
then dy/dx=0 will give you the point of inflexions
and d2y/dx2 will tell you about the concavity (i.e. concave up or down )the graph is.

BEST OF LUCK

We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL